OpenStudy (anonymous):

Write an equation in slope-intercept form for a line with m = −1/2 that passes through the point (2, 3).

OpenStudy (albert0898):

Slope intercept form is y = mx + b M, which is the slope is, -1/2 Plug the given coords in y = mx + b It would be 3 = -1/2(2) + b 3 = -1 + b Add 1 to both sides 4 = b I think you can figure the answer out using b = 4 and m = -1/2

OpenStudy (anonymous):

y=-1/2x+4